3. (a) Find the equivalent resistance of the circuit below with respect to point a and b.
(b) A battery of voltage ℇ = 8 V is connected in the circuit so that the point � to be the positive terminal of the battery and the point � the negative terminal. Calculate the current the battery produces to the circuit.
1
Expert's answer
2020-04-15T10:21:50-0400

we can see here there are 3 resistances connected in series to one another so the equivalent resistance will be equal to "R_1+R_2+R_3=8+4+3=15 ohm"

current in the circuit = "\frac VR=\frac{15}{15}= 1\ A"

now lets see the drop across each resistor

drop across R1 =8x1=8 volts

drop across R2 =4x1 =4volts

drop across R3 =3x1 =3 volts

current produces by the battery is 1 amperes

power delivered by the batttery = VI =15x1 =15 watt